OpenStudy (anonymous):

You deposit $1000 in an account that pays 6% annual intrest compounded continuously. find the balance at the end of 2 years

OpenStudy (anonymous):

in the case of compound interest use amount = principal * ((1+r/100)^n) here principal =1000 and r=6 and n=2 and u have to find amount hope it is easy for you to solve

OpenStudy (anonymous):

i got 1.0006

OpenStudy (anonymous):

no check your calculation again

OpenStudy (anonymous):

1.1236

OpenStudy (anonymous):

amount= 1000 *(1+6/100)^2 =1000*1.06*1.06=1000*1.1236=1123.6

OpenStudy (anonymous):

oh i forgot the *...thank you

OpenStudy (anonymous):

welcome dear